7.1x - 3.9y= 1.4




A temperature of 41 degrees F is equivalent to 278.2 degrees K, while a temperature of 86 degrees F is equivalent to 303.2 degrees K. Find the equation converting Kelvin temperature to Fahrenheit. What is the Fahrenheit temperature when it is 180 degrees K?

Answer by jojo14344(1513)

well, sorry there's no direct formula from deg F to kelvin.
But there's one for deg C to kelvin, so we need to convert it first to deg Celsius then to Kelvin (take note there's no "degree" in kelvin)
To show,
1.41 deg F to 278.2 kelvin
= (41-32)(5/9) = 5 deg C. in order to get in kelvin we need to add 273 (constant), well 273.15 if you want exact.
=5 + 273 = 278 kelvin
2. 86 deg F to 303.2 Kelvin
=(86-32)(5/9) = 30 deg C = 30 + 273 = 303 Kelvin
3. deg F? when 180Kelvin
we do it backwards,
=180- 273 = -93 deg C
For deg C to deg F,
= (-93 deg C)(9/5) + 32
=-135.4 deg F
cool!
Thank you,
jojo